OpenStudy (anonymous):

is this double chain rule? e^(xsin2x)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

double chain rule with a product rule

OpenStudy (anonymous):

which rule do i use? a^x'=a^xlna or e^u=e^udu/dx

OpenStudy (anonymous):

e^u=e^udu/dx

OpenStudy (anonymous):

[ e^u ] du u = xsin2x

OpenStudy (anonymous):

i get for du/dx x*cos2x*2+sin2x

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

cool beans